OpenStudy (anonymous):

What is the relationship between the lines determined by these two equations? 3x + 6y = 8 y = 2x – 8

OpenStudy (experimentx):

solve this, you will have solution, and at that solution, these lines meet each other

OpenStudy (hoblos):

slope of the first =m= -3/6=-1/2 slope of the first=m'=2 m*m' = (-1/2)*2 = -1 so the lines are perpendicular

OpenStudy (hoblos):

@experimentX !!!

OpenStudy (experimentx):

hahaha ... didn't note at all